f(x)=(1)/(√(4x^2-17x+4))
Find the domain and express in interval notation
info: its 1 over the square root of the numbers listed above

Answer by lwsshak3(11628)

f(x)=(1)/(√(4x^2-17x+4))
Find the domain and express in interval notation
info: its 1 over the square root of the numbers listed above
**
f(x)=(1)/(√(4x^2-17x+4))
=1/√(4x^2-17x+4)
=1/[√(4x-1)(x-4)]
check for values of x which makes function undefined:
4x-1=0
x≠1/4
or
x-4=0
x≠4
..
check for radican≥0
(4x-1)(x-4)≥0
number line
<..+....1/4....-......4.....+.........>
Domain: (-∞,1/4) U (4,∞)
In other words, all real numbers except x=1/4 and 4 and between x=1/4 and 4 (not including 1/4 and 4)